OpenStudy (anonymous):

What is the x-coordinate of the vertex of this parabola? y=-3x^2+12x-5

OpenStudy (anonymous):

\(-\frac{b}{2a}\)

OpenStudy (anonymous):

in your case \(a=-3,b=12\)

OpenStudy (anonymous):

A.)2 B.)-3 C.)12

OpenStudy (anonymous):

so which one

OpenStudy (anonymous):

which one is \(-\frac{b}{2a}=-\frac{12}{2\times (-3)}\) ?

OpenStudy (anonymous):

is A,B,or C